OpenStudy (anonymous):

ANSWERS GET MEDALS Find the distance from the center of the circle x^2+y^2-8x+6y=1 The point A(-5,-6)

myininaya (myininaya):

x^2-8x+y^2+6y=1 x^2-8x+4^2+y^2+6y+3^2=1+4^2+3^2 (x-4)^2+(y+3)^2=25 center of circle is (4,-3)

OpenStudy (anonymous):

centre is 4, -3 - then use pythagoras to find reqd distance

myininaya (myininaya):

d=sqrt{(-5-4)^2+(-6-(-3))^2} d=sqrt{81+9} d=sqrt{90}

OpenStudy (anonymous):

3sqrt10?